OpenStudy (anonymous):

$425 is 17% of what number?

OpenStudy (anonymous):

2500

OpenStudy (anonymous):

\[\frac{ is }{ of }=\frac{ percent }{ 100 }\] is a formula that makes this easy

OpenStudy (anonymous):

you would divide 425 by .17

OpenStudy (anonymous):

so its 2500 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[2500 \times 0.17=425\]

OpenStudy (anonymous):

not sure if i should put 2500.00 or 2500

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2500

OpenStudy (anonymous):

i put 2500 cuz it was 425 not 425.00

OpenStudy (anonymous):

Well, to answer this in more of an essay or short answer way would to do this.\[425\div0.17=2500\]This shows that you did the work.

OpenStudy (anonymous):

2500.00 is the same as 2500 though

OpenStudy (anonymous):

Technically, but if it's not in a dollar amount, then I would just put 2500

OpenStudy (anonymous):

Either way, I don't think that your teacher will care.

OpenStudy (anonymous):

yea ill just put 2500 if im wrong ill blame my self